OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f over g(−3) and select the correct answer below: f(x) = 6x + 8 g(x) = x − 2 −2 −negative one half one half 2

OpenStudy (anonymous):

\[\frac{ f }{ g }-3\]

OpenStudy (anonymous):

is it A or B?

OpenStudy (anonymous):

please, I want to work.. @mathway @Elsa213 @ganeshie8

OpenStudy (anonymous):

Make x=-3 in g(x). What's your answer?

OpenStudy (anonymous):

g(x) = x − 2 g(x)=-5

OpenStudy (anonymous):

Divide f(x) by g(x).

OpenStudy (anonymous):

f(x) = 6x + 8 f(x)=-10

OpenStudy (anonymous):

-10/-5, or -2

OpenStudy (anonymous):

i meant 2

OpenStudy (anonymous):

not negative

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

sorry my connection was lost :(

OpenStudy (anonymous):

its okay! :)

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

Substitute all x=-3.\[\frac{ 6x+8 }{ x-2}\]

OpenStudy (anonymous):

-10/-5

OpenStudy (anonymous):

or 2

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

Sorry ugh I just saw your last reply :( see, my connection is so bad today :(

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

haha its okay!Given the following functions f(x) and g(x), solve for (f ⋅ g)(2) and select the correct answer below: f(x) = 3x^2 + 2 g(x) = x − 8 −42 −84 42 84

OpenStudy (anonymous):

Or just substitute x=2 in f(x).

OpenStudy (anonymous):

then substitute x=2 in g(x)

OpenStudy (anonymous):

then multiply them together

OpenStudy (anonymous):

3*2^2 + 2*2 − 8 6^2+2*-6

OpenStudy (anonymous):

\[f(x)=3(2)^2+2\]\[g(x)=2-8\]

OpenStudy (anonymous):

14*-6

OpenStudy (anonymous):

Multiply them.

OpenStudy (anonymous):

-84, or B?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

(sorry caps)